# Courcelle’s Theorem for Lipschitz Continuity

### v1

@@ Abstract
keywords:
Fixed-Parameter Tractability, Algorithmic Meta-Theorem, Lipschitz Continuity
category:
[…]
@@ 1 Introduction
Lipschitz continuity of algorithms, introduced by Kumabe and Yoshida [KumabeY23], is a measure of an algorithm’s stability in response to errors or small perturbations in the input for weighted optimization problems. Roughly speaking, it is the maximum ratio of the (weighted) Hamming distance between the outputs of an algorithm for two different weights to the $\ell_{1}$ distance between those weights (see Section 2.4 for the precise definition). It is desirable for algorithms to have small Lipschitz constants, as large constants can undermine the reliability of decision-making and the reproducibility of scientific research.
Since its introduction, Lipschitz continuous algorithms have been proposed for various optimization problems [KumabeY23, KumabeY25]. However, we need to design a different algorithm and analyze its Lipschitz continuity for each problem, which can be impractical. To address this limitation, we present an algorithmic meta-theorem for Lipschitz continuous algorithms, which can be seen as Lipschitz continuity analogue of celebrated Courcelle’s theorem [Courcelle90].
@@
We note that the trivial upper bound on Lipschitz constant is $n$ [KumabeY25]; therefore the bounds in the theorems above are significantly smaller for fixed $\mathrm{tw}$ and $\varphi$. We also remark that our meta-theorems yield randomized approximation algorithms. This is necessary since, for most problems, it is known that exact or deterministic algorithms cannot be Lipschitz continuous [KumabeY23].
When the treewidth of the input graph is bounded by a constant, Theorems 1.1 and 1.2 provide algorithms that outperform existing algorithms in terms of approximability and/or Lipschitz continuity:
  • For the minimum weight vertex cover problem, Theorem 1.2 yields an algorithm with a better approximation ratio than the previous $2$-approximation algorithm with Lipschitz constant $4$ [KumabeY25], at the polylogarithmic sacrifice of the Lipschitz constant.
  • For the minimum weight feedback vertex set problem, Theorem 1.2 outperforms the previous $O(\log n)$-approximation algorithm with Lipschitz constant $O(\sqrt{n}\log^{3/2}n)$ [KumabeY25] in terms of both approximability and Lipschitz continuity.
  • For the maximum weight matching problem [^1] Although the output of this problem is an edge set rather than a vertex set, this problem can be expressed by an $\textsf{MSO}_{2}$ formula on a new graph $G^{\prime}=(V\cup E,\{(v,e)\colon e\text{ is incident to }v\})$, Theorem 1.1 yields an algorithm with a better approximation ratio than the previous $\left(\frac{1}{8}-\varepsilon\right)$-approximation algorithm with Lipschitz constant $O(\varepsilon^{-1})$ [KumabeY23], at the polylogarithmic sacrifice of the Lipschitz constant.
  • For the shortest path problem, Theorem 1.2 slightly improves the Lipschitz continuity compared to the previous $(1-\varepsilon)$-approximation algorithm with Lipschitz constant $O(\varepsilon^{-1}\log^{3}n)$ [KumabeY23], without losing approximability.
For a fixed $\varphi$, by explicitly specifying the transitions in the dynamic programming within the algorithm, we can provide more precise bounds on the function $f$ that appears in Theorems 1.1 and 1.2. In particular, considering the case that $\varphi$ is an $\textsf{MSO}_{2}$ formula representing the independent set constraint, we have the following.
Theorem 1.3.
@@
We further demonstrate that Theorems 1.1 and 1.2 lead to Lipschitz continuous version of Baker’s technique [Baker94]. As a representative example, we consider the maximum weight independent set problem on planar graphs, where a vertex subset $X$ is an independent set of a graph $G$ if no two vertices in $X$ are adjacent in $G$. We prove the following.
Theorem 1.4.
@@
As a lower bound result, we prove that the Lipschitz constant of any $(1-\varepsilon)$-approximation algorithm for some $\textsf{MSO}_{2}$ maximization problem is large on general graphs, which justifies considering the $\textsf{MSO}_{2}$ maximization problem on a restricted graph class. Specifically, we consider the max ones problem [creignou2001complexity, khanna2001approximability], where the instance is a 3CNF formula over a variable set $X$ and a weight function $w:X\to\mathbb{R}_{\geq 0}$, and the goal is to find a satisfying assignment $\sigma:X\to\{0,1\}$ that maximizes the weight $\sum_{x\in\sigma^{-1}(1)}w(x)$. It is easy to see that there exists a fixed $\textsf{MSO}_{2}$ formula $\varphi(X)$ such that the max ones problem can be reduced to a problem on bipartite graphs, where the task is to find a vertex set $X$ that maximizes the weight $\sum_{x\in X}w(x)$ subject to $\varphi$. We show the following.
Theorem 1.5.
@@
We note that it is possible to construct an algorithm for the max ones problem with a Lipschitz constant similar to that in Theorem 1.3, where $\mathrm{tw}$ denotes the treewidth of the “incidence graph” of the 3CNF instance (see Section 4.2 for details). Since $\mathrm{tw}$ is upper bounded by $n$ for any instance, Theorem 1.5 implies that the dependency of the Lipschitz constant on $\mathrm{tw}$ cannot be improved to $\mathrm{tw}^{o(1)}$.
@@
We note that the functions $f$ in Theorems 1.6 and 1.7 are much larger than those in Theorems 1.1 and 1.2. In particular, the bounds for Theorems 1.3 and 1.4 do not follow from Theorem 1.6.
@@ 1.1 Technical Overview
However, this algorithm is not Lipschitz continuous. This is because when we compute $\mathsf{DP}[v][1]$, which of $X_{v,0}$ or $X_{v,1}$ has a larger weight is affected by small changes in the weights. To address this issue, we use the exponential mechanism [Dwork06]. Specifically, for some constant $c>0$, we select $X_{v,i}$ with probability proportional to $\exp(c\cdot w(X_{v,i}))$. This approach makes the algorithm Lipschitz continuous, with only a slight sacrifice in the approximation ratio. Specifically, we can prove that by appropriately choosing $c$ for a given $\varepsilon^{\prime}\in(0,1]$, the increase in the Lipschitz constant can be bounded by $\tilde{O}(\varepsilon^{\prime-1})$ by reducing the approximation guarantee by a factor of $1-\varepsilon^{\prime}$ at each vertex where the exponential mechanism is applied.
While this approach makes the algorithm Lipschitz continuous, the Lipschitz constant is still too large when the height $h$ of the tree is large. Specifically, to achieve an approximation guarantee of $1-\varepsilon$, we need to set $\varepsilon^{\prime}=\frac{\varepsilon}{h}$, leading to a Lipschitz constant of $h\cdot\tilde{O}(\varepsilon^{\prime-1})=\tilde{O}(\varepsilon^{-1}h^{2})$. This is larger than the trivial bound $n$ when $h=O(n)$. We resolve this issue by using the fact that any tree has a tree decomposition of width 5 and height $O(\log n)$ [BodlaenderH98]. By performing dynamic programming with the exponential mechanism on this tree decomposition, we can obtain $(1-\varepsilon)$-approximation algorithm with Lipschitz constant $\tilde{O}(\varepsilon^{-1})$. This argument can be naturally extended to the case where $G$ is a bounded treewidth graph. By following the proof of Courcelle’s Theorem [Courcelle90, CourcelleM93], we further extend this argument to the case where $\varphi$ is a general $\textsf{MSO}_{2}$ formula with a free vertex variable.
We prove Theorem 1.5 by leveraging a lower bound on a related notion of sensitivity for the maximum cut problem [fleming2024sensitivity], where sensitivity measures the change in the output with respect to the Hamming distance under edge deletions [varma2023average]. Specifically, we reduce the maximum cut problem, where stability is defined with respect to edge deletions, to the max ones problem, where stability is measured with respect to weight changes.
@@ 1.2 Related Work
Lipschitz continuity of discrete algorithms is defined by Kumabe and Yoshida [KumabeY23], and they provided Lipschitz continuous approximation algorithms for the minimum spanning tree, shortest path, and maximum weight matching problems. In a subsequent work [KumabeY25], they further provided such algorithms for the minimum weight vertex cover, minimum weight set cover, and minimum weight feedback vertex set problems. In another work [KumabeY24], they defined Lipschitz continuity for allocations in cooperative games and provided Lipschitz continuous allocation schemes for the matching game and the minimum spanning tree game.
A variant known as pointwise Lipschitz continuity has also been studied, which is defined using the unweighted Hamming distance instead of the weighted Hamming distance. Kumabe and Yoshida [KumabeY23] defined pointwise Lipschitz continuity and provided pointwise Lipschitz continuous algorithms for the minimum spanning tree and maximum weight bipartite matching problems. Liu et al. [liu2024pointwise] proposed the proximal gradient method as a general technique for solving LP relaxations stably. Using this, they provided pointwise Lipschitz continuous approximation algorithms for the minimum vertex $(S,T)$-cut, densest subgraph, maximum weight ($b$-)matching, and packing integer programming problems.
(Average) sensitivity, introduced by Varma and Yoshida [varma2023average] with preliminary ideas by Murai and Yoshida [murai2019sensitivity], is a notion similar to Lipschitz continuity. While Lipschitz continuity evaluates an algorithm’s stability against unit changes in the input weights, (average) sensitivity evaluates an algorithm’s stability against (random) deletions of elements from the input. Algorithms with small (average) sensitivity have been studied for several problems, such as maximum matching [varma2023average, YoshidaZ21], minimum cut [varma2023average], knapsack problem [KumabeY22], Euclidean $k$-means [yoshida2022average], spectral clustering [peng2020average], and dynamic programming problems [kumabe2022average]. Recently, Fleming and Yoshida [fleming2024sensitivity] constructed a PCP framework to prove the sensitivity lower bound for the constraint satisfaction problem.
@@ 1.3 Organization
The rest of this paper is organized as follows. In Section 2, we describe the necessary preliminaries on tree decomposition (Sections 2.1 and 2.2), logic (Section 2.3), and Lipschitz continuity (Section 2.4). In Section 3, we prove Theorem 1.1 by providing a Lipschitz continuous algorithm for the $\textsf{MSO}_{2}$ maximization problem. Since the algorithm and analysis for $\textsf{MSO}_{2}$ minimization are similar, we defer the proof of Theorem 1.2 to Appendix C. In Section 4, we give a more precise Lipschitzness analysis for specific formulas $\varphi$, which includes the proof of Theorem 1.3. In Section 5, we prove Theorem 1.4 by providing a Lipschitz continuous version of Baker’s technique. In Section 6, we discuss a lower bound on the Lipschitz constant for general graphs, which justifies focusing on graphs with bounded treewidth. In Section 7, we prove Theorem 1.6 by providing a Lipschitz continuous algorithm for the $\textsf{MSO}_{1}$ maximization problem, while we defer the proof of Theorem 1.7 to Appendix C.
@@ 2.1 Tree Decomposition
For a tree decomposition $(\mathcal{B},\mathcal{T})$, we may refer to the root node, leaf nodes, and the height of $\mathcal{T}$ as those of $(\mathcal{B},\mathcal{T})$, respectively. Moreover, $(\mathcal{B},\mathcal{T})$ is binary if $\mathcal{T}$ is a binary tree. The width of a tree decomposition $(\mathcal{B},\mathcal{T})$ is the maximum size of a bag in $\mathcal{B}$ minus one. The treewidth of $G$ is the minimum possible width among all possible tree decompositions of $G$. It is known that the tree decomposition of $G$ of width at most $12k+$ 1 can be computed in $2^{O(k)}n$ time [Korhonen21], where $k$ is the treewidth of $G$. Moreover, any tree decomposition of $G$ of width $k$ can be transformed into a binary tree decomposition of width at most $23k+$ 2 and height at most $O(\log n)$ [BodlaenderH98]. Thus, we obtain the following lemma.
Lemma 2.1 ([BodlaenderH98, Korhonen21]).
@@ 2.2 HR-algebra
We introduce the notions of HR-algebra [^2] The acronym HR stands for hyperedge replacement [CourcelleE2012]., which is one of the algebraic definitions of tree decomposition. A $k$-graph is a tuple $G=\langle V,E,\mathsf{src}\rangle$ of a set of vertices $V$, a set of edges $E$, and a $k$-vector $\mathsf{src}\in(V\cup\{\mathtt{nil}\})^{k}$. We write the $i$-th element of a vector $\mathsf{src}$ by $\mathsf{src}(i)$. If $\mathsf{src}(i)=u$, we say that $u$ is a $i$-source. We write the set $\{\mathsf{src}(i):i\in[k]\}\setminus\{\mathtt{nil}\}$ by $\mathsf{src}(G)$.
@@
The proof of Proposition 2.4 is essentially the same as the proof of Proposition 2.3 but we give in Appendix A.
@@ 2.3 Monadic Second-Order Logic
We provide a slightly less formal definition of monadic second-order (MSO) logic over $k$-graphs for accessibility (see, e.g., [Libkin04] for more formal definition). An atomic formula is a formula of the form $s=t$, $\mathsf{adj}(s,t)$, $t\in X$, and Boolean constants $\mathtt{True}$ and $\mathtt{False}$. Here, $\mathsf{adj}(s,t)$ is the predicate that represents whether $s$ and $t$ are adjacent. A monadic second-order formula over $k$-graphs is a formula built from atomic formulas, the usual Boolean connectives ,,¬$\land$,$\lor$,$\lnot$,$\to$ , ∨ , ¬ , →, first-order quantifications $\exists x$ and $\forall x$ (quantifications over vertices), and second-order quantifications $\exists X$ and $\forall X$ (quantification over sets of vertices). The notation $\varphi(X)$ indicates that $\varphi$ has exactly one free variable $X$. The quantifier height (or quantifier rank) of a formula $\varphi$, denoted by $|\varphi|$, is the maximum depth of quantifier nesting in $\varphi$ (see e.g. [Libkin04] for a formal definition). The counting monadic second-order (CMSO) logic is an expansion of MSO logic that have additional predicates $\mathsf{Card}_{m,r}(X)$ for a second-order variable $X$, meaning $|X|\equiv m\pmod{r}$. A $\textsf{C}_{r}\textsf{MSO}^{q}$-formula is a CMSO formula $\varphi$ such that the quantifier height of $\varphi$ is at most $q$ and $r^{\prime}\leq r$ holds for any predicate $\mathsf{Card}_{m,r^{\prime}}$ in $\varphi$. For a logical formula $\varphi$ and a graph $G$, we write $G\models\varphi$ when graph $G$ satisfies the property expressed by $\varphi$.
There are two variants of MSO logic: $\textsf{MSO}_{1}$ (or simply MSO), which is described above, allows quantification over vertices and sets of vertices only, and $\textsf{MSO}_{2}$ which allows quantification over edges and sets of edges as well. It is well known that, for any graph $G$ and $\textsf{MSO}_{2}$-formula $\varphi$, there exists an $\textsf{MSO}_{1}$ formula $\varphi^{\prime}$ such that $G\models\varphi\iff G^{\prime}\models\varphi^{\prime}$, where $G^{\prime}$ is the incidence graph, obtained by subdividing each edge of $G$ (with two colors meaning the original and the subdivision vertices) (see e.g. [Kreutzer11]). It is easy to see that the treewidth of the incidence graph is at most the treewidth of the original graph. Thus, since this paper focuses on graphs with bounded treewidth, we mainly consider $\textsf{MSO}_{1}$. However, all the theorems are held for $\textsf{MSO}_{2}$ and graphs of bounded treewidth.
@@
The following theorem is a key of the MSO model-checking algorithm [Courcelle90, CourcelleE2012].
Theorem 2.5 ([CourcelleE2012]).
@@
It is known that, for any $r,q,k$, up to logical equivalence, there are only finitely many different $\textsf{C}_{r}\textsf{MSO}^{q}$-formulas $\varphi(X)$ over $k$-graphs (see e.g., [Libkin04]). Thus, we can obtain a linear-time algorithm, based on dynamic programming over trees, for $\textsf{C}_{r}\textsf{MSO}^{q}$ model checking over bounded treewidth. Courcelle and Mosbah [CourcelleM93] adjusted Theorem 2.5 to address optimization, counting, and other problems. We use a slightly modified version of the theorem of Courcelle and Mosbah.
Let $\mathsf{nosrc}(X)\equiv\bigwedge_{i\in[k]}(\mathsf{src}(i)\not\in X)$, and $\varphi\mathord{{\restriction}_{S}}(X)$ denote $\varphi(X\cup S)\land\mathsf{nosrc}(X)$ for a set $S\subseteq\mathsf{src}(G)$. {observation} Let $G$ be a $k$-graph, and $\varphi(X)$ be a formula with a free set variable $X$. Then, we have $\mathsf{sat}(G,\varphi)=\biguplus_{S\subseteq\mathsf{src}(G)}\mathsf{sat}(G,\varphi\mathord{{\restriction}_{S}})\boxtimes\{S\}$.
The proof of Corollary 2.6 is given in Appendix A
Corollary 2.6 ($\star$).

### v2

@@ 1 Introduction
Lipschitz continuity of algorithms, introduced by Kumabe and Yoshida [15], is a measure of an algorithm’s stability in response to errors or small perturbations in the input for weighted optimization problems. Roughly speaking, it is the maximum ratio of the (weighted) Hamming distance between the outputs of an algorithm for two different weights to the $\ell_{1}$ distance between those weights (see Section 2.4 for the precise definition). It is desirable for algorithms to have small Lipschitz constants, as large constants can undermine the reliability of decision-making and the reproducibility of scientific research.
Since its introduction, Lipschitz continuous algorithms have been proposed for various optimization problems [15, 17]. However, we need to design a different algorithm and analyze its Lipschitz continuity for each problem, which can be impractical. To address this limitation, we present an algorithmic meta-theorem for Lipschitz continuous algorithms, which can be seen as Lipschitz continuity analogue of celebrated Courcelle’s theorem [3].
@@
We note that the trivial upper bound on Lipschitz constant is $n$ [17]; therefore the bounds in the theorems above are significantly smaller for fixed $\mathrm{tw}$ and $\varphi$. We also remark that our meta-theorems yield randomized approximation algorithms. This is necessary since, for most problems, it is known that exact or deterministic algorithms cannot be Lipschitz continuous [15].
When the treewidth of the input graph is bounded by a constant, Theorems 1 and 2 provide algorithms that outperform existing algorithms in terms of approximability and/or Lipschitz continuity:
  • For the minimum weight vertex cover problem, Theorem 2 yields an algorithm with a better approximation ratio than the previous $2$-approximation algorithm with Lipschitz constant $4$ [17], at the polylogarithmic sacrifice of the Lipschitz constant.
  • For the minimum weight feedback vertex set problem, Theorem 2 outperforms the previous $O(\log n)$-approximation algorithm with Lipschitz constant $O(\sqrt{n}\log^{3/2}n)$ [17] in terms of both approximability and Lipschitz continuity.
  • For the maximum weight matching problem [^1] Although the output of this problem is an edge set rather than a vertex set, this problem can be expressed by an $\textsf{MSO}_{2}$ formula on a new graph $G^{\prime}=(V\cup E,\{(v,e)\colon e\text{ is incident to }v\})$, Theorem 1 yields an algorithm with a better approximation ratio than the previous $\left(\frac{1}{8}-\varepsilon\right)$-approximation algorithm with Lipschitz constant $O(\varepsilon^{-1})$ [15], at the polylogarithmic sacrifice of the Lipschitz constant.
  • For the shortest path problem, Theorem 2 slightly improves the Lipschitz continuity compared to the previous $(1-\varepsilon)$-approximation algorithm with Lipschitz constant $O(\varepsilon^{-1}\log^{3}n)$ [15], without losing approximability.
For a fixed $\varphi$, by explicitly specifying the transitions in the dynamic programming within the algorithm, we can provide more precise bounds on the function $f$ that appears in Theorems 1 and 2. In particular, considering the case that $\varphi$ is an $\textsf{MSO}_{2}$ formula representing the independent set constraint, we have the following.
Theorem 3.
@@
We further demonstrate that Theorems 1 and 2 lead to Lipschitz continuous version of Baker’s technique [1]. As a representative example, we consider the maximum weight independent set problem on planar graphs, where a vertex subset $X$ is an independent set of a graph $G$ if no two vertices in $X$ are adjacent in $G$. We prove the following.
Theorem 4.
@@
As a lower bound result, we prove that the Lipschitz constant of any $(1-\varepsilon)$-approximation algorithm for some $\textsf{MSO}_{2}$ maximization problem is large on general graphs, which justifies considering the $\textsf{MSO}_{2}$ maximization problem on a restricted graph class. Specifically, we consider the max ones problem [7, 10], where the instance is a 3CNF formula over a variable set $X$ and a weight function $w:X\to\mathbb{R}_{\geq 0}$, and the goal is to find a satisfying assignment $\sigma:X\to\{0,1\}$ that maximizes the weight $\sum_{x\in\sigma^{-1}(1)}w(x)$. It is easy to see that there exists a fixed $\textsf{MSO}_{2}$ formula $\varphi(X)$ such that the max ones problem can be reduced to a problem on bipartite graphs, where the task is to find a vertex set $X$ that maximizes the weight $\sum_{x\in X}w(x)$ subject to $\varphi$. We show the following.
Theorem 5.
@@
We note that it is possible to construct an algorithm for the max ones problem with a Lipschitz constant similar to that in Theorem 3, where $\mathrm{tw}$ denotes the treewidth of the “incidence graph” of the 3CNF instance (see Section 4.2 for details). Since $\mathrm{tw}$ is upper bounded by $n$ for any instance, Theorem 5 implies that the dependency of the Lipschitz constant on $\mathrm{tw}$ cannot be improved to $\mathrm{tw}^{o(1)}$.
@@
We note that the functions $f$ in Theorems 6 and 7 are much larger than those in Theorems 1 and 2. In particular, the bounds for Theorems 3 and 4 do not follow from Theorem 6.
@@ 1.1 Technical Overview
However, this algorithm is not Lipschitz continuous. This is because when we compute $\mathsf{DP}[v][1]$, which of $X_{v,0}$ or $X_{v,1}$ has a larger weight is affected by small changes in the weights. To address this issue, we use the exponential mechanism [8]. Specifically, for some constant $c>0$, we select $X_{v,i}$ with probability proportional to $\exp(c\cdot w(X_{v,i}))$. This approach makes the algorithm Lipschitz continuous, with only a slight sacrifice in the approximation ratio. Specifically, we can prove that by appropriately choosing $c$ for a given $\varepsilon^{\prime}\in(0,1]$, the increase in the Lipschitz constant can be bounded by $\tilde{O}(\varepsilon^{\prime-1})$ by reducing the approximation guarantee by a factor of $1-\varepsilon^{\prime}$ at each vertex where the exponential mechanism is applied.
While this approach makes the algorithm Lipschitz continuous, the Lipschitz constant is still too large when the height $h$ of the tree is large. Specifically, to achieve an approximation guarantee of $1-\varepsilon$, we need to set $\varepsilon^{\prime}=\frac{\varepsilon}{h}$, leading to a Lipschitz constant of $h\cdot\tilde{O}(\varepsilon^{\prime-1})=\tilde{O}(\varepsilon^{-1}h^{2})$. This is larger than the trivial bound $n$ when $h=O(n)$. We resolve this issue by using the fact that any tree has a tree decomposition of width 5 and height $O(\log n)$ [2]. By performing dynamic programming with the exponential mechanism on this tree decomposition, we can obtain $(1-\varepsilon)$-approximation algorithm with Lipschitz constant $\tilde{O}(\varepsilon^{-1})$. This argument can be naturally extended to the case where $G$ is a bounded treewidth graph. By following the proof of Courcelle’s Theorem [3, 6], we further extend this argument to the case where $\varphi$ is a general $\textsf{MSO}_{2}$ formula with a free vertex variable.
We prove Theorem 5 by leveraging a lower bound on a related notion of sensitivity for the maximum cut problem [9], where sensitivity measures the change in the output with respect to the Hamming distance under edge deletions [24]. Specifically, we reduce the maximum cut problem, where stability is defined with respect to edge deletions, to the max ones problem, where stability is measured with respect to weight changes.
@@ 1.2 Related Work
Lipschitz continuity of discrete algorithms is defined by Kumabe and Yoshida [15], and they provided Lipschitz continuous approximation algorithms for the minimum spanning tree, shortest path, and maximum weight matching problems. In a subsequent work [17], they further provided such algorithms for the minimum weight vertex cover, minimum weight set cover, and minimum weight feedback vertex set problems. In another work [16], they defined Lipschitz continuity for allocations in cooperative games and provided Lipschitz continuous allocation schemes for the matching game and the minimum spanning tree game.
A variant known as pointwise Lipschitz continuity has also been studied, which is defined using the unweighted Hamming distance instead of the weighted Hamming distance. Kumabe and Yoshida [15] defined pointwise Lipschitz continuity and provided pointwise Lipschitz continuous algorithms for the minimum spanning tree and maximum weight bipartite matching problems. Liu et al. [19] proposed the proximal gradient method as a general technique for solving LP relaxations stably. Using this, they provided pointwise Lipschitz continuous approximation algorithms for the minimum vertex $(S,T)$-cut, densest subgraph, maximum weight ($b$-)matching, and packing integer programming problems.
(Average) sensitivity, introduced by Varma and Yoshida [24] with preliminary ideas by Murai and Yoshida [20], is a notion similar to Lipschitz continuity. While Lipschitz continuity evaluates an algorithm’s stability against unit changes in the input weights, (average) sensitivity evaluates an algorithm’s stability against (random) deletions of elements from the input. Algorithms with small (average) sensitivity have been studied for several problems, such as maximum matching [24, 26], minimum cut [24], knapsack problem [14], Euclidean $k$-means [25], spectral clustering [22], and dynamic programming problems [13]. Recently, Fleming and Yoshida [9] constructed a PCP framework to prove the sensitivity lower bound for the constraint satisfaction problem.
@@ 1.3 Organization
The rest of this paper is organized as follows. In Section 2, we describe the necessary preliminaries on tree decomposition (Sections 2.1 and 2.2), logic (Section 2.3), and Lipschitz continuity (Section 2.4). In Section 3, we prove Theorem 1 by providing a Lipschitz continuous algorithm for the $\textsf{MSO}_{2}$ maximization problem. Since the algorithm and analysis for $\textsf{MSO}_{2}$ minimization are similar, we defer the proof of Theorem 2 to Appendix C. In Section 4, we give a more precise Lipschitzness analysis for specific formulas $\varphi$, which includes the proof of Theorem 3. In Section 5, we prove Theorem 4 by providing a Lipschitz continuous version of Baker’s technique. In Section 6, we discuss a lower bound on the Lipschitz constant for general graphs, which justifies focusing on graphs with bounded treewidth. In Section 7, we prove Theorem 6 by providing a Lipschitz continuous algorithm for the $\textsf{MSO}_{1}$ maximization problem, while we defer the proof of Theorem 7 to Appendix C.
@@ 2.1 Tree Decomposition
For a tree decomposition $(\mathcal{B},\mathcal{T})$, we may refer to the root node, leaf nodes, and the height of $\mathcal{T}$ as those of $(\mathcal{B},\mathcal{T})$, respectively. Moreover, $(\mathcal{B},\mathcal{T})$ is binary if $\mathcal{T}$ is a binary tree. The width of a tree decomposition $(\mathcal{B},\mathcal{T})$ is the maximum size of a bag in $\mathcal{B}$ minus one. The treewidth of $G$ is the minimum possible width among all possible tree decompositions of $G$. It is known that the tree decomposition of $G$ of width at most $12k+$ 1 can be computed in $2^{O(k)}n$ time [11], where $k$ is the treewidth of $G$. Moreover, any tree decomposition of $G$ of width $k$ can be transformed into a binary tree decomposition of width at most $23k+$ 2 and height at most $O(\log n)$ [2]. Thus, we obtain the following lemma.
Lemma 8 ([2, 11]).
@@ 2.2 HR-algebra
We introduce the notions of HR-algebra [^2] The acronym HR stands for hyperedge replacement [4]., which is one of the algebraic definitions of tree decomposition. A $k$-graph is a tuple $G=\langle V,E,\mathsf{src}\rangle$ of a set of vertices $V$, a set of edges $E$, and a $k$-vector $\mathsf{src}\in(V\cup\{\mathtt{nil}\})^{k}$. We write the $i$-th element of a vector $\mathsf{src}$ by $\mathsf{src}(i)$. If $\mathsf{src}(i)=u$, we say that $u$ is a $i$-source. We write the set $\{\mathsf{src}(i):i\in[k]\}\setminus\{\mathtt{nil}\}$ by $\mathsf{src}(G)$.
@@
The proof of Proposition 11 is essentially the same as the proof of Proposition 10 but we give in Appendix A.
@@ 2.3 Monadic Second-Order Logic
We provide a slightly less formal definition of monadic second-order (MSO) logic over $k$-graphs for accessibility (see, e.g., [18] for more formal definition). An atomic formula is a formula of the form $s=t$, $\mathsf{adj}(s,t)$, $t\in X$, and Boolean constants $\mathtt{True}$ and $\mathtt{False}$. Here, $\mathsf{adj}(s,t)$ is the predicate that represents whether $s$ and $t$ are adjacent. A monadic second-order formula over $k$-graphs is a formula built from atomic formulas, the usual Boolean connectives ,,¬$\land$,$\lor$,$\lnot$,$\to$ , ∨ , ¬ , →, first-order quantifications $\exists x$ and $\forall x$ (quantifications over vertices), and second-order quantifications $\exists X$ and $\forall X$ (quantification over sets of vertices). The notation $\varphi(X)$ indicates that $\varphi$ has exactly one free variable $X$. The quantifier height (or quantifier rank) of a formula $\varphi$, denoted by $|\varphi|$, is the maximum depth of quantifier nesting in $\varphi$ (see e.g. [18] for a formal definition). The counting monadic second-order (CMSO) logic is an expansion of MSO logic that have additional predicates $\mathsf{Card}_{m,r}(X)$ for a second-order variable $X$, meaning $|X|\equiv m\pmod{r}$. A $\textsf{C}_{r}\textsf{MSO}^{q}$-formula is a CMSO formula $\varphi$ such that the quantifier height of $\varphi$ is at most $q$ and $r^{\prime}\leq r$ holds for any predicate $\mathsf{Card}_{m,r^{\prime}}$ in $\varphi$. For a logical formula $\varphi$ and a graph $G$, we write $G\models\varphi$ when graph $G$ satisfies the property expressed by $\varphi$.
There are two variants of MSO logic: $\textsf{MSO}_{1}$ (or simply MSO), which is described above, allows quantification over vertices and sets of vertices only, and $\textsf{MSO}_{2}$ which allows quantification over edges and sets of edges as well. It is well known that, for any graph $G$ and $\textsf{MSO}_{2}$-formula $\varphi$, there exists an $\textsf{MSO}_{1}$ formula $\varphi^{\prime}$ such that $G\models\varphi\iff G^{\prime}\models\varphi^{\prime}$, where $G^{\prime}$ is the incidence graph, obtained by subdividing each edge of $G$ (with two colors meaning the original and the subdivision vertices) (see e.g. [12]). It is easy to see that the treewidth of the incidence graph is at most the treewidth of the original graph. Thus, since this paper focuses on graphs with bounded treewidth, we mainly consider $\textsf{MSO}_{1}$. However, all the theorems are held for $\textsf{MSO}_{2}$ and graphs of bounded treewidth.
@@
The following theorem is a key of the MSO model-checking algorithm [3, 4].
Theorem 12 ([4]).
@@
It is known that, for any $r,q,k$, up to logical equivalence, there are only finitely many different $\textsf{C}_{r}\textsf{MSO}^{q}$-formulas $\varphi(X)$ over $k$-graphs (see e.g., [18]). Thus, we can obtain a linear-time algorithm, based on dynamic programming over trees, for $\textsf{C}_{r}\textsf{MSO}^{q}$ model checking over bounded treewidth. Courcelle and Mosbah [6] adjusted Theorem 12 to address optimization, counting, and other problems. We use a slightly modified version of the theorem of Courcelle and Mosbah.
Let $\mathsf{nosrc}(X)\equiv\bigwedge_{i\in[k]}(\mathsf{src}(i)\not\in X)$, and $\varphi\mathord{{\restriction}_{S}}(X)$ denote $\varphi(X\cup S)\land\mathsf{nosrc}(X)$ for a set $S\subseteq\mathsf{src}(G)$.
Observation 13.
Let $G$ be a $k$-graph, and $\varphi(X)$ be a formula with a free set variable $X$. Then, we have $\mathsf{sat}(G,\varphi)=\biguplus_{S\subseteq\mathsf{src}(G)}\mathsf{sat}(G,\varphi\mathord{{\restriction}_{S}})\boxtimes\{S\}$.
The proof of Corollary 14 is given in Appendix A
Corollary 14 ($\star$).